simplify without using a calculator:
1.) 8^2/3
2.) 16^-3/2
3.) numerator is 2^1/5
denominator is 2^-6/5

I really Appreciate this! thank you so much!

Answer by ewatrrr(24785)


Hi
simplify without using a calculator:
1.) 8^2/3 (cubed root of 8, squared)= 4
2.) 16^-3/2 (1/square root of 16,cubed) = 1/64
3.) (2^1/5 )/ (2^-6/5) is 2^7/5